------------------------------------------------------------ Date: Sun Jan 20 15:47:12 2002, Shift: 08:00-16:00 Supervisor: R. Debbe, with: Log book#12 pages 94 - 95 Shift Summary: ------------------------------------------------------------ 1. There was a store and data taking when I arrived, continued taking data till 10:17, after that, the machine was highjacked by masked polarized people and it looks like they are still holding their ground. 2. Dana was here in the morning, we tried to budge BFS but failed. He is trying to contact the engineers to see if we can do some work during access tomorrow. 3. BFS is stuck at 3 degrees, I was unable to submit my shift check. ------------------------------------------------------------
This archive was generated by hypermail 2b30 : Sun Jan 20 2002 - 15:47:49 EST